### v1

# Generative Artificial Intelligence for Beamforming in Low-Altitude Economy

Abstract
The growth of low-altitude economy (LAE) has driven a rising demand for efficient and secure communication. However, conventional beamforming optimization techniques struggle in the complex LAE environments. In this context, generative artificial intelligence (GenAI) methods provide a promising solution. In this article, we first introduce the core concepts of LAE and the roles of beamforming in advanced communication technologies for LAE. We then examine their interrelation, followed by an analysis of the limitations of conventional beamforming methods. Next, we provide an overview of how GenAI methods enhance the process of beamforming, with a focus on its applications in LAE. Furthermore, we present a case study using a generative diffusion model (GDM)-based algorithm to enhance the performance of aerial collaborative beamforming-enabled remote secure communications in LAE and simulation results verified the effectiveness of the proposed algorithms. Finally, promising research opportunities are identified.
Index Terms:
generative artificial intelligence, beamforming, low-altitude economy, generative diffusion model
I Introduction
Low-altitude economy (LAE) is an economic model that relies on the utilization of low-altitude airspace for various commercial activities, with widespread applications in the fields such as precision agriculture, urban logistics, and disaster rescue [1]. By enabling efficient aerial resource allocation, LAE can quickly adapt to the demands of high-density users and dynamic tasks. Moreover, LAE enhances the service response speed and coverage by enabling rapid resource deployment, especially in remote or congested areas where traditional infrastructure may be limited or unavailable. However, to meet these demands, LAE faces challenging requirements on communication systems, to support efficient and flexible resource management in rapidly changing environments [2].
As a core technology in wireless communication systems, beamforming has been widely applied in LAE networking. By controlling the phase and amplitude of antenna arrays, beamforming enhances signal directionality and strength, thereby addressing the challenges such as signal coverage, interference suppression, and spectrum allocation. However, conventional beamforming optimization methods face limitations in LAE scenarios, which involve dynamic channel conditions, high-density interference, and complex task requirements. These methods struggle to adapt to real-time environmental changes and lack sufficient flexibility for multi-user collaboration and resource allocation. Therefore, relying solely on conventional beamforming optimization methods is insufficient to fully meet the communication demands of LAE networking.
Fortunately, recent advances in artificial intelligence (AI) technologies offer promising solutions to these issues. Traditional AI (TAI) techniques primarily target discriminative tasks such as classification and regression, where the main objective is to map inputs to predefined outputs. In the context of beamforming optimization, TAI techniques typically map observed channel conditions directly to predefined optimal beamforming parameters. Although these TAI methods perform well for classification or regression tasks in known scenarios, they inherently depend on labeled datasets and fixed optimization objectives. In contrast, generative artificial intelligence (GenAI) models focus on identifying the underlying distribution of the data. By capturing the intrinsic data distributions, GenAI models establish probabilistic models that identify hidden relationships within unstructured environmental data. Consequently, GenAI methods are capable of synthesizing optimal beamforming configurations through probabilistic inference, thereby generalizing beyond historical scenarios. This shift from mapping inputs to predefined outputs towards a data-driven generative approach fundamentally improves the adaptability of dynamic networks and simultaneously reduces the reliance on labeled datasets.
Figure 1: This figure presents the two key technologies in beamforming: beam management and beam tracking, each of which illustrates how it contributes to optimizing communication performance.
By applying generative models, GenAI methods can capture complex channel characteristics and generate highly accurate channel state information (CSI), thereby improving the design and adjustment strategies for beamforming. Moreover, GenAI models present potential for adapting to frequent channel variations in dynamic environments and enhancing resource utilization and interference management. Additionally, GenAI models can learn from diverse data distributions, which enables them to evolve given various communication requirements and lays the foundation for multi-scenario applications within LAE.
Motivated by the potential of GenAI models in addressing these challenges, we focus on integrating them with beamforming within the LAE. The main contributions of this work are as follows:
  • We first introduce the characteristics of LAE and the fundamental concepts of beamforming. Then, we elaborate on the applications of beamforming in both communications and sensing in the context of LAE. Finally, we discuss the challenges faced by beamforming in LAE scenarios.
  • We describe several core techniques and applications of GenAI. Subsequently, we discuss how these GenAI techniques can be applied to beam management and beam tracking, as well as their specific roles in beamforming optimization for LAE.
  • We propose a framework based on GenAI to optimize the performance of aerial collaborative beamforming-enabled remote secure communications in LAE. Simulation results show that the proposed framework significantly improves the secrecy rate and energy consumption of the considered system.
II Overview of Low-altitude Economy Beamforming
In this section, we first introduce the core concepts of LAE. Next, we outline the fundamental principles and processes of beamforming. Finally, we explore the various applications of beamforming in LAE and emphasize its roles in communications and sensing.
II-A Low-Altitude Economy
LAE is an emerging economic model that uses low-altitude airspace to tackle the challenges of modern cities. By incorporating advanced communication technologies and intelligent computing into low-altitude platforms (LAPs), LAE improves the efficiency and responsiveness of industries such as logistics, urban mobility, and emergency services. The key characteristics and challenges of LAE are as follows.
II-A1 Diverse Mission Types
LAE is marked by diverse missions. For instance, LAPs are used for aerial photography, and help with property promotion. Also, in medical emergencies, LAPs transport crucial supplies. Thus, this variety calls for customized configurations and specialized operations.
II-A2 Decentralized Communication
LAE uses ad hoc communications, where aerial vehicles communicate directly with each other without a centralized controller. This enables flexible coordination and ensures reliability in environments where traditional infrastructure may be limited or unavailable [3].
II-A3 Complex Environments
Low-altitude operations are affected by urban structures, weather conditions, and obstacles. Therefore, adaptive sensing and communication strategies are essential to ensure stable and reliable operations.
Given the aforementioned characteristics and challenges of LAE, effective wireless signal management and optimization in low-altitude operations are crucial. Hence, beamforming emerges as a relevant solution, which has substantial potential in tackling signal-related challenges.
Figure 2: Beamforming enables efficient communications and sensing applications in the LAE, such as interference avoidance, spectrum reuse, target tracking, and environmental monitoring.
II-B Beamforming
Beamforming is a signal processing technique for optimizing the directionality of electromagnetic waves by controlling the phase and amplitude of each antenna element in the array antenna. Fig. 1 illustrating the key technologies involved in beamforming is provided, and the details of these technologies are as follows.
II-B1 Beam Management
Beam management focuses on acquiring and maintaining reliable beam pairs to ensure robust directional communication. In LAE, it maintains stable links for certain applications such as drone logistics and aerial monitoring, which ensures uninterrupted communication for mission-critical tasks. To achieve this, beam management includes four key steps, which work together to maintain optimal communication paths in real time. These steps are as follows.
  • Beam Sweeping: Beam sweeping involves scanning predefined beam directions to identify optimal signal paths. In LAE, rapid beam sweeping addresses the high mobility of LAPs, thus enabling quick alignment with ground nodes or other aerial devices.
  • Beam Measurement: Beam measurement evaluates signal quality metrics including strength and interference across candidate beams. In LAE, this process dynamically adapts to environmental fluctuations such as weather changes or urban reflections to sustain robust link quality.
  • Beam Determination: Based on the results of beam measurements, beam determination selects the optimal beam direction and width. This step is critical for LAE to prioritize low-latency links for time-sensitive tasks, such as collision avoidance or real-time surveillance.
  • Beam Reporting: Beam reporting involves transmitting feedback on beam status to network controllers or centralized systems. This process supports coordinated multi-LAP operations, thereby minimizing interference and optimizing spectrum utilization.
II-B2 Beam Tracking
Beam tracking dynamically adjusts the beam direction in real time to follow the movement of targets. In LAE operations, where targets such as moving vehicles or other LAPs frequently change their positions, beam tracking helps LAPs maintain stable communication links.
II-C Applications of Beamforming in Low-Altitude Economy
Beamforming technology supports a wide range of applications in low-altitude networking, broadly classified into communication and sensing domains. Specifically, Fig. 2 highlights the classical applications within these domains. Meanwhile, a comprehensive literature review diagram is presented in Fig. 3, which offers an overview of the low-altitude networking, key beamforming techniques, and their applications in low-altitude networking.
II-C1 Communications
Beamforming plays a critical role in LAE communications. First, beamforming reduces interference by directing signal energy more accurately toward a specified target through the enhanced directivity of the antenna array. This advantage becomes especially significant in dense urban environments and large-scale events, where multiple signal sources compete for limited spectrum resources [4]. With another aspect of beam focusing, beamforming strengthens security by making signal interception significantly more challenging for unintended parties. Moreover, beamforming enhances spectrum utilization via effective interference mitigation by tailoring beams considering both the direction and distance according to the location of each user or device, which enables more efficient spectrum reuse and allows multiple users to share the same frequency band simultaneously.
II-C2 Sensing
Beamforming can effectively address the sensing tasks in LAE applications. First, beamforming enhances the ability of LAPs to track targets by adjusting the beam direction based on the real-time position of the target, which ensures consistent focus and maintains tracking accuracy even with changing trajectories. Furthermore, beamforming strengthens the ability of LAPs to monitor the environment. Specifically, by focusing wireless signal energy, beamforming increases signal strength in target areas, which helps resolve attenuation during long-distance transmission [5].
Figure 3: An overview of the timeline presenting the developments in low-altitude networking, beamforming techniques, and applications of beamforming in low-altitude networking.
II-D Existing Challenges
The effectiveness of conventional beamforming technologies is currently constrained by several unresolved challenges, which are further discussed as follows.
II-D1 Scalability Constraints
As wireless communication systems continue to evolve, the scale of deployment in terms of antenna arrays, user density, and spatial coverage is rapidly increasing. In LAE scenarios, where multiple platforms may operate simultaneously, beamforming techniques must be capable of handling large and complex configurations. However, conventional approaches often rely on fixed model assumptions and manually designed optimization procedures, which become increasingly difficult to maintain as the system scales. As a result, this lack of scalability limits their effectiveness in dense and dynamic communication scenarios.
II-D2 Real-Time Adaptation
Although current adaptive beamforming methods can handle certain dynamic adjustments, they often struggle to achieve real-time adaptation in LAE. Tasks such as LAP delivery and low-altitude mapping involve rapidly changing flight paths and wireless channel conditions, which require immediate updates to beam directions. However, conventional approaches rely on infrequent update cycles and limited feedback mechanisms, which lack the responsiveness to track rapid variations in CSI under highly dynamic conditions.
II-D3 Multi-Objective Optimization
Beamforming balances the interference reduction, energy efficiency, and communication reliability. However, it often faces stringent environmental and resource constraints. As a result, overlapping paths and limited spectrum in LAE scenarios make emergency communications and similar high-demand situations increasingly challenging. Therefore, addressing these multi-objective challenges calls for more advanced optimization techniques.
II-D4 Antenna Technology Complexity
Conventional beamforming techniques have significant limitations when faced with advanced antenna technologies like massive multiple-input multiple-output (MIMO), holographic MIMO, and extremely-large MIMO. The reason is that conventional beamforming designs rely on mathematical and analytical models and struggle to handle the complexity and high demands of these advanced technologies. As a result, conventional beamforming designs cause lower efficiency and limited adaptability, thus highlighting the need for alternative technologies.
The aforementioned challenges highlight the need for more intelligent and adaptable solutions to improve beamforming in LAE applications. In this context, GenAI technologies provide flexible and efficient methods to address the complexity and dynamic requirements of modern beamforming.
TABLE I: The Use of Generative AI in Beamforming and Beamforming in Low-Altitude Economy.
Issues Generative AI Model Potential Benefits
GAN VAE GDM Transformer
Beam Management Simulate robust feedback channels Denoise multimodal inputs Optimize beam direction through probabilistic generation and refinement Identify relationships across modalities [6] 1) Improve the accuracy of beamforming 2) Increase the real-time responsiveness of beamforming 3) Reduce the computational overhead of beamforming
Beam Tracking Generate diverse environmental training datasets [7] Encode channel features into latent space Enable parallel tracking with multi-head attention
Beamforming in LAE Improve channel modeling and estimation [8] [9] and reduce energy Extract features to generate beam and spectrum allocation strategies Generate adaptive beam directions through learned patterns Capture long-range dependencies and model context [10] 1) Enhance channel characterization and estimation 2) Optimize energy consumption, resource allocation, and security
III Generative AI for Beamforming in Low-Altitude Economy
In this section, we begin by introducing foundational GenAI models. Next, we discuss how GenAI enhances beamforming and highlight specific advancements in the context of LAE, as presented in Table I.
III-A Fundamentals of Generative AI
III-A1 Generative Adversarial Network (GAN)
GANs are a class of GenAI models that consist of two competing neural networks, which are widely used in areas such as image generation and data augmentation. The core principle of GANs lies in adversarial training between the generator and the discriminator, where the generator progressively enhances the realism of the generated data while the discriminator improves its ability to distinguish real data from fake. In wireless communications, GANs contribute to mitigating noise by modeling complex data distributions, thereby leading to more stable and robust communication. For instance, in [11], a GAN-based model was designed to improve bit error rate (BER) performance in uplink Rayleigh fading channels. As a result, the decreased BER improves the accuracy of channel state information (CSI) estimation. This then allows for better adjustment of beamforming weights to focus energy on target users while suppressing interference, which increases the signal-to-interference-plus-noise ratio (SINR). Moreover, the enhanced SINR facilitates more accurate CSI estimation through cleaner feedback. During this process, beamforming achieves precise signal alignment by dynamically adjusting phase and amplitude, performing matched filtering through the channel to maximize array gain, ultimately leading to improved overall communication performance.
III-A2 Variational Autoencoder (VAE)
Apart from GANs, VAEs represent another major class of GenAI models that learn latent representations of input data and generate similar outputs by maximizing the data likelihood. Each model consists of an encoder that compresses the data into latent parameters, and a decoder that reconstructs the data from this distribution. In wireless communications, the unique encoding and decoding mechanisms of VAEs enhance information security. For example, in [12], the authors proposed a VAE-based authentication scheme that encodes channel impulse response features into a probabilistic latent space, enabling secure authentication without requiring the channel prior knowledge of attacker. Thus, VAEs can support secure beamforming by encoding channel features into latent distributions, dynamically optimizing parameters to resist interference.
III-A3 Generative Diffusion Model (GDM)
In addition to GANs and VAEs, GDMs have emerged as a powerful class of GenAI models based on probabilistic denoising processes, which generate data through forward and reverse diffusion processes. Specifically, the forward process gradually adds noise to transform the data into a Gaussian distribution, while the reverse process removes the noise to reconstruct the data and recover the target distribution. In wireless communications, GDMs are valuable for channel simulation, as they can replicate complex channel characteristics. For example, in [13], the authors applied a GDM-based model to generate channel data conditioned on user positions, producing realistic MIMO channel samples even with limited real measurements. Consequently, beamforming algorithms can leverage these simulated channel conditions to improve beam direction estimation, particularly in sparse channel measurement scenarios.
III-A4 Transformer
Transformer models are deep learning architectures centered around attention mechanisms and have gained prominence in various generative tasks. In this paradigm, input data representing temporal and spatial variations in network behavior is first converted into tokens. Unlike GANs, VAEs, and GDMs, which generate samples by sampling from a learned latent distribution, Transformer-based models generate content by modeling the conditional probability of each element in a sequence, producing outputs in either an autoregressive or parallel manner. In wireless communications, Transformer models can facilitate channel estimation tasks. For example, in [14], the authors embedded dynamic network states into structured tokens and employed a Transformer to capture multi-scale temporal and spatial dependencies. The resulting representations effectively modeled time-varying channel dynamics and user preferences, thereby enhancing the exploration performance of actor-critic networks in deep reinforcement learning. As such, Transformer-based approaches can optimize beamforming by dynamically adjusting beam parameters in response to real-time network conditions.
III-B Generative AI for Beamforming
III-B1 Generative AI Solutions in Beam Management
Conventional beam management approaches face notable limitations in dynamic and complex environments. First, conventional beam management approaches employ oversimplified beam prediction models that lack the adaptability to rapid channel variations caused by mobility and environmental obstructions. Additionally, these methods primarily utilize single-modality data while ignoring multimodal contextual information from supplementary sensing inputs like visual perception systems. Furthermore, rigid spatial coordination strategies and linear optimization frameworks fail to capture the long-range spatiotemporal correlations, resulting in suboptimal beam alignment and compromised throughput efficiency. In contrast, GenAI-enhanced methods address these limitations by leveraging multi-modal data fusion and adaptive learning.
For instance, in [6], the authors proposed a multi-modal Transformer model combined with reinforcement learning to optimize beam management. Specifically, the model processes visual, light detection and ranging (LiDAR), and radar data by using self-attention mechanisms to identify the relationships across modalities. This enables efficient prediction of optimal beam groups by reducing the computational complexity through narrowed search spaces.
Beyond holistic beam management, GenAI techniques can further enhance specific sub-processes. For beam sweeping, GDMs could synthesize diverse environmental scenarios such as varying pedestrian densities or weather conditions to optimize initial beam scans. In beam measurement, VAEs can denoise LiDAR or radar inputs to enhance signal quality and ensure reliable beam measurements. Furthermore, the feature fusion capability of multi-modal Transformer directly aids beam determination by integrating multi-modal data to prioritize high-SINR beams. Finally, for beam reporting, GANs could simulate realistic feedback channels, thus ensuring robust reporting under latency or packet loss.
III-B2 Generative AI Solutions in Beam Tracking
Conventional beam tracking algorithms require real-time calculations of the movement of target devices and channel states to adjust the beam direction. However, in complex environments with interference and blockages, rapid changes in channel conditions can make it difficult for tracking to stay aligned with the target. Additionally, conventional methods often rely on simple motion models, such as linear or Kalman filter-based predictions, which struggle to accurately predict the highly uncertain movement trajectories of user devices. In contrast, GenAI models can generate comprehensive training datasets covering various environmental scenarios and complex movement trajectories.
For instance, in [7], the authors proposed a GAN-based method to generate diverse and comprehensive training datasets. These datasets include variations such as different human postures, environmental changes, as well as multipath and blockage scenarios. Furthermore, such datasets can be used to enhance the training of deep learning models, thereby improving their ability to adapt to rapidly changing environments and ensuring more stable and accurate beam tracking.
III-C Generative AI for Beamforming in Low-Altitude Economy
GenAI elevates beamforming in LAE by dynamically adapting to complex environmental uncertainties and collaborative network demands. This section demonstrates its capabilities in mission-critical tasks and resilient performance enhancement.
III-C1 Communication Tasks
The mobility and interference of LAE demand agile beamforming. GenAI transcends conventional approaches, thus enabling robust task execution, as detailed in two key tasks.
  • Channel Modeling and Estimation: Compared to conventional communication scenarios, LAE places higher demands on channel modeling and estimation due to its complex and highly dynamic environment. Specifically, factors such as adverse weather conditions, obstruction by buildings or terrain features, interference from surrounding aerial and ground-based nodes, and the rapid mobility of LAPs significantly impact signal quality. These external factors often lead to severe signal attenuation and pronounced multipath propagation effects, thus posing substantial challenges to accurate channel modeling and estimation. Although conventional beamforming techniques can adapt to certain channel variations, they fall short in effectively addressing the rapid temporal and spatial fluctuations introduced by these external environmental factors. Therefore, there is a clear need for innovative approaches, and GenAI offers promising solutions by leveraging advanced data-driven frameworks to improve channel modeling and estimation.
    For instance, Zhang et al. [8] developed a GAN-based model to improve channel estimation in millimeter-wave communications for low-altitude unmanned aerial vehicle (UAV) networks. Specifically, the model allows each UAV to independently train channel models for specific beamforming directions, with samples shared through a collaborative framework to build a generalized model, which improves estimation accuracy. Additionally, Rasheed et al. [9] proposed a distributed GAN model based on long short-term networks for low-altitude maritime UAV communications. Specifically, UAVs utilize the model to extract real-time CSI and generate synthetic samples, which improve the accuracy of channel state estimation.
  • Resource Allocation: Compared to conventional communication environments, efficient spectrum and power resource allocation in LAE scenarios is significantly more challenging due to dense node deployments, rapid network topology changes caused by node mobility, and intense competition for limited resources. However, conventional beamforming approaches often fail to quickly adjust beam configurations in response to these complex factors, thus leading to inefficient resource management. In contrast, GenAI-based beamforming methods explicitly link beamforming optimization with resource allocation, enabling simultaneous and real-time adjustment of beam directions and spectrum assignments to address resource conflicts and enhance utilization efficiency.
    For instance, a VAE-based framework can encode network topology variations and resource competition conditions into concise latent features. Subsequently, the decoder directly generates optimized beamforming parameters and adaptive spectrum allocation schemes. This integrated approach effectively mitigates interference, resolves resource competition conflicts, and improves overall spectrum and power utilization efficiency in highly dynamic LAE scenarios.
Figure 4: The workflow of the GDMTD3 algorithm includes environmental observation, action generation via a diffusion-based actor network, evaluation by twin critic networks, and policy updates supported by a replay buffer.
III-C2 Performance Metrics
LAE demands adaptive security and energy resilience against evolving threats. GenAI delivers fundamental breakthroughs, as detailed in the following two aspects.
  • Communication Security: Compared to the conventional communication scenarios, communication security in LAE is particularly critical. On the one hand, different tasks may rely on different networks or frequency bands, which means that the LAPs frequently need to switch between multiple communication networks and frequency bands. As a result, this increases the vulnerability of communication links to network attacks or malicious interference. On the other hand, LAE includes tasks in critical areas such as logistics and emergency rescue, where the reliability of communication security directly impacts the efficiency and safety of these services. However, conventional beamforming algorithms rely solely on signal strength to optimize beam direction, which limits their ability to handle dynamic malicious interference. In contrast, GenAI models excel at analyzing latent data distributions and dynamically optimizing anti-jamming strategies under uncertain conditions.
    For example, in [10], the authors combined a diffusion-based optimization framework with a mixture of expert (MoE)-Transformer network, and leveraged multi-head attention to model spatiotemporal channel dependencies. Moreover, this method dynamically optimizes the three-dimensional beamforming to resist eavesdropping interference, mitigate location errors, and address channel uncertainties, thus enhancing the anti-jamming resilience and secrecy performance in the mission-critical low-altitude communication systems.
  • Energy Consumption: In LAE senarios, the frequent beam redirection by LAPs, interference suppression between high-density nodes, and multi-network handovers require continuous power consumption, thus leading to a significant increase in energy usage compared to conventional communication environments. However, conventional beamforming algorithms rely on pre-set power distribution strategies, which hinder the ability to balance communication quality and energy efficiency in dynamic topology and channel fluctuation scenarios. GenAI can learn the relationship between environmental features and energy consumption online, autonomously generating beam control strategies that optimize energy efficiency and overcome the limitations of traditional static approaches.
    For instance, a beam control system based on GAN can employ a generator to simulate low-power beam patterns and dynamically optimize beam width and radiation power. Meanwhile, the discriminator can assess signal integrity based on real-time channel states, which ensures a reduction in energy consumption without sacrificing performance.
III-D Lessons Learned
In this section, we explored the application of GenAI models in beamforming, particularly within the context of LAE. From these insights, we derive the following key lessons.
  1. 1.
    GenAI models fundamentally enhance beamforming adaptability in dynamic environments through real-time parameter generation and high-fidelity channel modeling. This capability directly addresses the rapid spatial-temporal variations and interference challenges specific to low-altitude scenarios while maintaining robust performance under mobility constraints and environmental uncertainties.
  2. 2.
    By compressing high-dimensional channel data into latent representations and generating synthetic training samples, GenAI models reduce computational overhead while simultaneously enhancing resource allocation, beam tracking, and energy efficiency. This dual capability of optimizing performance while minimizing complexity is crucial for real-time operations in resource-constrained LAE scenarios.
  3. 3.
    The ability of GenAI models to to model evolving threats and environmental dynamics supports proactive security. By synthesizing attack patterns and channel anomalies, GenAI models can generate adaptive anti-jamming strategies while maintaining reliable beamforming performance, which is essential for mission-critical LAE applications such as emergency communications.
IV Case Study: Generative AI for aerial collaborative beamforming in Low-Altitude Economy
In this section, we examine a GenAI-based approach for collaborative UAV beamforming in LAE scenarios. First, we discuss the motivation behind this approach and provide an overview of the system design. Next, we introduce the proposed algorithm. Finally, we evaluate how this method enhances the communication efficiency and security.
IV-A Motivation and System Design
In LAE scenarios, security risks are rising due to complex terrains and high-density LAP deployments, which often lead to vulnerabilities in communication, with traditional encryption and anti-jamming methods struggling under resource constraints [15]. Additionally, energy efficiency is critical for LAPs due to their limited power supply and high energy consumption. To address these challenges, this case study utilizes UAVs, key components of LAPs, to implement collaborative aerial beamforming for secure remote communications. Specifically, multiple UAVs form a virtual antenna array to enhance the signal directionality and strength towards the remote base station, while optimizing excitation current weights and UAV positions to ensure secure and energy-efficient data transmission. The proposed optimization framework combines the GDM model with DRL to improve both security and energy efficiency.
IV-B Proposed Algorithm
To address the challenges identified in the system design, this study proposes the GDM-enhanced twin delayed deep deterministic policy gradient (GDMTD3) algorithm, which integrates a GDM into the twin delayed deep deterministic policy gradient (TD3) framework. This integration allows the algorithm to optimize UAV swarm actions by modeling the problem as a Markov decision process. Specifically, the state is defined by the positions of the UAVs and the estimated location of the eavesdropper. Based on this, the action consists of selecting excitation current weights and adjusting UAV positions. Furthermore, the reward function combines the secrecy rate and energy consumption, while also imposing penalties for violations such as exceeding speed limits or UAV collisions.
Fig. 4 shows the framework of the proposed GDMTD3 algorithm. Specifically, the workflow begins with the system observing the current environment, which includes the positions of the UAVs and the estimated location of the eavesdropper. Based on this information, the diffusion-based actor network processes the state and generates the candidate actions. Specifically, GDM replaces the standard actor network in the TD3 framework, where it learns the probability distribution of optimal actions and applies a reverse diffusion process. This allows the algorithm to refine noisy inputs into high-quality actions that are better tailored to the current environment. Subsequently, these actions are evaluated by the twin critic networks, which compute the corresponding secrecy rate and energy consumption. In turn, the evaluations provide crucial feedback that guides the actor policy, thus helping it make optimal trade-offs between security and efficiency. Furthermore, the replay buffer stores historical state-action-reward tuples, which are then sampled to iteratively update both the actor and critic networks. This iterative process ensures policy convergence to actions that enhance communication security, minimize energy consumption, and maintain stability in training. Through this closed-loop process, the algorithm dynamically adapts to environmental changes, thereby achieving robust performance in both security and energy efficiency.
Figure 5: Performance evaluation of the GDMTD3 algorithm compared with DDPG, SAC, and TD3. (a) Average reward per episode. (b) Average secrecy rate and energy consumption per step.
IV-C Performance Evaluation
To validate the effectiveness of the proposed GDMTD3 algorithm, extensive simulations were conducted and compared with baseline algorithms, including deep deterministic policy gradient (DDPG), soft actor-critic (SAC), and standard TD3. The experimental findings are summarized as follows.
As shown in Fig. 5(a), the GDMTD3 algorithm achieved higher and more stable rewards across 1000 training iterations compared to the baseline methods. This demonstrates its ability to learn optimal policies effectively and adapt to dynamic environments, while the baseline algorithms exhibited slower convergence and greater variability.
In terms of optimization objectives, Fig. 5(b) shows that the proposed GDMTD3 achieved the highest secrecy rate with the lowest energy consumption per step, effectively balancing the security and efficiency. In contrast, SAC and TD3 showed moderate performance in one metric but failed to achieve the optimal trade-off.
V Conclusion and Future Directions
In this article, we have introduced the review of applying GenAI models in beamforming for LAE. We have first analyzed the communication requirements in LAE and the limitations of conventional beamforming techniques. Next, we have explored the applications of GenAI models in two aspects, namely beam management and beam tracking, and highlighted its potential to enhance the communication performance. Subsequently, we have presented a case study to verify the effectiveness of the proposed GDM-based collaborative beamforming method in improving the communication security and energy efficiency for LAE. In addition, three other directions for future research in GenAI techniques for beamforming in LAE scenarios are outlined below.
V-A Semantic-Driven Beamforming
As communication demands in LAE scenarios shift from traditional data transmission to semantic information delivery, existing beamforming techniques struggle to meet precise requirements. In response, integrating beamforming with semantic communication frameworks is a promising direction. This approach dynamically adjusts beam directions and power allocation based on semantic task demands. As a result, it can better support evolving communication needs.
V-B Long-Range High-Power Wireless Power Transfer-Driven Beamforming
Unlike short-range wireless power transfer (WPT) targeting low-power devices and sensors, LAPs require long-range, high-power WPT for reliable and sustainable operation. In this context, it is necessary to realize long-range high-power WPT with beamforming, such as laser and RF-based power beamforming. Moreover, integrating space-based solar power with such advanced WPT solutions can address energy challenges, ultimately supporting the development of 6G systems across various applications in LAE.
V-C Integrated Sensing and Communications-Driven Beamforming
Multiple independent sensing beams are necessary to detect separate targets in integrated sensing and communications (ISAC) scenarios anticipated for 6G systems, particularly in LAE applications. This is due to the fact that near-field beam focusing with large-scale MIMO technology concentrates transmit signal energy in specific directions and distances, which differs from the far-field ISAC systems. Consequently, future research on ISAC-driven beamforming (i.e., beam focusing) for LAE applications should prioritize the use of multiple narrow beams to optimize sensing performance and ensure efficient operation of LAPs.